OpenStudy (anonymous):

Anyone wanna medal? If g(x) = x2 + 2, find g(3). A. 9 B. 8 C. 11 D. 6

OpenStudy (jadeishere):

All you have to do with this is plug in the value of 3 into the 'x' so g(x) = 3^2 + 2 3^2 = 9 so g(x) = 9 + 2 then you'll get your answer >.< I hope this helps :)

OpenStudy (anonymous):

Thank's

OpenStudy (anonymous):

Can you help me with a similar question?

OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. 12 2.6 8 4

OpenStudy (jadeishere):

Do what I said, just plug in the value >.<

OpenStudy (anonymous):

g(x)= 3(6-2) ?

OpenStudy (anonymous):

Should it be 12, A?

OpenStudy (jadeishere):

Yep :) Now if you solve that equation, you'll get the answer >.< PEMDAS so solve the parentheses first 3 * (6-2) is your answer >.<

OpenStudy (anonymous):

claps